Please provide a list of all trignometric formulas in class 11!

Grade:11

3 Answers

Manish
6 years ago

List of Trigonometry Class 11 Formulas

sin(θ)=sinθ
cos(θ)=cosθ
tan(θ)=tanθ
cosec(θ)=cosecθ
sec(θ)=secθ
cot(θ)=cotθ

Product to Sum Formulas

sinx siny=12[cos(xy)cos(x+y)]
cosxcosy=12[cos(xy)+cos(x+y)]
sinxcosy=12[sin(x+y)+sin(xy)]
cosxsiny=12[sin(x+y)sin(xy)]

Sum to Product Formulas

sinx+siny=2sin(x+y2)cos(xy2)
sinxsiny=2cos(x+y2)sin(xy2)
cosx+cosy=2cos(x+y2)cos(xy2)
cosxcosy=2sin(x+y2)sin(xy2)

Basic Formulas

sin(A+B)=sinAcosB+cosAsinB sin(AB)=sinAcosBcosAsinB cos(A+B)=cosAcosBsinAsinB cos(AB)=cosAcosB+sinAsinB tan(A+B)=tanA+tanB1tanAtanB tan(AB)=tanAtanB1+tanAtanB cos(A+B)cos(AB)=cos2Asin2B=cos2Bsin2A sin(A+B)sin(AB)=sin2Asin2B=cos2Bcos2A sin2A=2sinAcosA=2tanA1+tan2A cos2A=cosAsin2A=12sin2A=2cos2A1=1tan2A1+tan2A tan2A=2tanA1tan2A
Please excuse me that I couldnt put the divided by lines wherever there is ann empty space assume lines there.
Hope it helped you!
ankit yadav
6 years ago
Salo apni maa chudaonahi to apni behan chuda lekoi mil nahi raha to mujhe bula lebehanchod madarchod
ankit yadav
6 years ago
sorry mere dost ne majak majak mein likh diya tha dont mindmaf kar do bhaisorry gjgtwgt.hjtgqh.hjuwdtagj.gthph